Project

General

Profile

Actions

Physics Shift Report

Start Date: 12-10-2017
End Date: 12-14-2017
Start Run: 10474
End Run: 10514
Physics shifter: Tanaz A. Mohayai

Summary of Runs

Run Number Run Type Absorber Momentum Cooling Channel Fields
10474 - 10477 Pulser Run LiH 240 Field Off
10477 Pulser Run LiH 240 Flip - 2T, 2017-02-7
10478 Special Run LiH 140 Flip - 2T, 2017-02-7
10479 Reference Run LiH N/A Flip - 2T, 2017-02-7
10480 - 10490 Special Run LiH 140 Flip - 2T, 2017-02-7
10491, 10492 Pulser Run LiH 140, 170 Flip - 2T, 2017-02-7
10493 - 10502 Special Run LiH 170 Flip - 2T - 2017-02-8a
10503 Pulser Run LiH 170 Field Off
10504 - 10511 Special Run LiH 240 Solenoid - 2T - 2017-02-6
10512 - 10513 Test - Beamline Tuning LiH 140 Solenoid - 2T - 2017-02-6
10514 Pulser Run - DAQ flush LiH N/A Field Off

Elog, CDB, and Reconstruction

Elog entries are ok.

CDB uploaded ok with the exception of the following runs:
10475
Note that runs 10491-10492 labeled as Pulser runs are checks on the datamover
Online reconstruction is ok
All runs are reconstructing ok

Updated by Mohayai, Tanaz Angelina over 4 years ago · 12 revisions